fuel cells lectricity. 18.51 Calculate the amounts of Cu and Br, produced in 1.0 h at inert electrodes in a solution of CuBr 18.51
Answer #1

Q = I x t; I is current in amp ; t is time in seconds

    = 4.5 x 1 x 60 x 60

    = 16200 C

Cu2+ + 2e- \rightarrow Cu

2F required for 1 mole of Cu = 63.5 g/mol

Mass of copper deposited at cathode = (63.5 /2 x 96500 C ) x 16200 C= 5.330 g

Similarly,   2 Br-   \rightarrow    Br2 +  2e-

2F required for 1 mole of Br = 159.5 g/mol

Mass of bromine deposited at anode = (159.5/2 x 96500 C) x 16200 C = 13.363 g

Note: Please give your feedback